# Judging from Support-set: A New Way to Utilize Few-Shot Segmentation for Segmentation Refinement Process

### v2

@@ Introduction
Table 1: Comparison of the mIoU of the coarse mask, the SEPL refined mask, and JFS picked segmentations, along with the success rate of JFS across different data selections. The success rate was measured as the number of correct decisions out of the total samples. Top 50/20 refers to the 50/20 samples with the greatest improvement in the refinement process, while Low 50/20 denotes the 50/20 samples with the least improvement. The best results are highlighted in bold.
@@
Quantitative results: Table 1 compares the mIoU for coarse, refine and JFS picked segmentations, and shows the success rate of JFS across different experiment groups. The results show that the JFS picked segmentations consistently outperform the SEPL refinement in all four groups. The highest improvement by JFS was observed in the low 20 samples (where SEPL struggles), where JFS improved mIoU to 50.5 from SEPL’s 33.2. In this group, both methods showed lower mIoU than the original coarse masks, but JFS improves on SEPL outputs by detecting failure cases in SEPL outputs. Furthermore, the success rate, which measures the effectiveness of JFS, is higher where mIoU is significantly improved by SEPL.
Qualitative results: Figure 2 and Figure 4 show that JFS can determine failure and success cases in the segmentation refinement process. Therefore, if the segmentation refinement fails, we can return to the existing coarse mask and consider other segmentation refinement methods. For this visualization, we used images publicly available on Wikimedia Commons for support-set. Query images are from PASCAL. We selected images where only one object was visible for each class. The support mask was obtained using the SAM segmentation method.
@@
How JFS works?: Most FSS networks (Shaban et al. 2017; Zhang et al. 2019; Tian et al. 2020; Liu et al. 2020; Wang et al. 2020; Min, Kang, and Cho 2021; Hong et al. 2022; Moon et al. 2022, 2023) rely on high quality support set information to find target objects in the query image (Figure 1). Therefore, FSS performance deteriorates when the support mask information is inaccurate (Moon et al. 2022, 2023). If a refined mask accurately segments the target object, it provides more detailed texture information of the target to FSS model. Hence, we designed JFS. We found from our qualitative analysis (Figure 2 and Figure 4) that the JFS method was particularly effective in the following cases :
@@
(b) When the difference between the coarse and refined mask is clear, the presence of non-target objects in the refined mask confuses the FSS model, significantly dropping support mask prediction accuracy. Thus, it becomes evident whether the refined mask has failed or succeeded. Table 1 shows the JFS success rate rises with significant mIoU improvement, indicating JFS’s higher effectiveness when refinement module makes greater improvement.
@@
Limitations of JFS: Table 1 suggests that the success rate of JFS decreases when there is a minor improvement in the refinement process. Specifically, our observations reveal the limitations of JFS in two cases (Figure 5).
(a) If the refined mask is not significantly different from the existing segmentation, JFS struggles as $E^{c}$ and $E^{r}$ values are similar. Therefore, JFS performs better in deciding on cases where the refined mask shows a significant difference from the coarse mask.
(b) When both the refined mask and the coarse mask share the key feature of the object, FSS network will generate similar support mask predictions and the results of JFS need to be carefully considered in such cases.
Figure 5: Failure cases. When the refined mask and coarse mask cover the same key features on the same target object, support mask prediction yields similar results. This is because FSS models are trained to find a target even with small key features from support-set. JFS failed to decide accurately whether the refinement process was successful or failed.
Path to Deployment
Addressing the limitations outlined above is crucial to improving the performance of JFS. There are two potential strategies to improve the performance of JFS. The first approach is to develop and integrate a more effective segmentation refinement tool. A robust refinement tool could significantly enhance the precision and reliability of the initial segmentation output, laying a stronger foundation for subsequent processes. The second approach is to deploy a more accurate few-shot segmentation (FSS) model in the second stage of JFS. The advanced FSS model would not only improve the overall segmentation performance, but also increase the likelihood of correctly identifying successful refinements. These approaches can substantially improve the success rate of JFS. As advancements in AI continue, particularly in the realm of few-shot segmentation models, there is still room for improving our framework with newly developed segmentation models. The continuous evolution of these models will enable more nuanced and precise segmentation, which can be leveraged within the JFS framework to achieve superior results.
In addition, there is also an opportunity to incorporate language descriptions as a novel form of guidance in evaluating refinement success as vision-language models (VLM) become increasingly sophisticated. Our framework has the potential to integrate semantic understanding into the refinement process by harnessing the descriptive capabilities of VLM, allowing more context-aware and accurate assessments. This potential integration of linguistic and visual information could lead to a more holistic and effective segmentation refinement strategy, extending the boundaries of what can be achieved with JFS. Generative AI is also a promising tool that can support the JFS framework by generating similar images and masks based on the query image, which can then be used as support masks. The aforementioned approaches would enable JFS to achieve higher accuracy in detecting even fine-detail differences.
Conclusion
We proposed JFS, a novel method for assessing the effectiveness of the segmentation refinement process. In this paper, we demonstrate that JFS has the capability to distinguish between successful and failed cases in SEPL segmentation refinement. The identification of successful refinements can be applied in various scenarios, leading to improved segmentation quality. Furthermore, JFS introduces a fresh perspective in the domains of few-shot segmentation and segmentation refinement by repurposing FSS for a unique application. However, we recognize that JFS underperforms in certain instances, indicating the need for further development. Future research could achieve better results by fine-tuning FSS models or adapting them specifically for the JFS framework. Moreover, exploring the integration of additional modalities, such as linguistic information, or leveraging generative AI could provide new avenues to enhance the JFS method. This continued exploration and refinement of JFS will contribute to expanding the boundaries of segmentation refinement and enriching its practical applications.

### v3

@@ 4 Result and Discussion
Table 1: Comparison of mIoU of the coarse mask, the SEPL-refined mask, and JFS-selected segmentations, along with the success rate of JFS across different data selections. The success rate is measured as the percentage of correct decisions out of total samples. ”Top 50/20” refers to the 50/20 samples with the greatest improvement in the refinement process, while ”Low 50/20” denotes the 50/20 samples with the least improvement. The best results are highlighted in bold.
@@
Quantitative results: Table 1 compares the mIoU for coarse, refine and JFS picked segmentations, and shows the success rate of JFS across different experiment groups. The results show that the JFS picked segmentations consistently outperform the SEPL refinement in all four groups. The highest improvement by JFS was observed in the low 20 samples (where SEPL struggles), where JFS improved mIoU to 50.5 from SEPL’s 33.2. In this group, both methods showed lower mIoU than the original coarse masks, but JFS improves on SEPL outputs by detecting failure cases in SEPL outputs. Furthermore, the success rate, which measures the effectiveness of JFS, is higher where mIoU is significantly improved by SEPL. Qualitative results: Fig. 2 shows that JFS can determine failure and success cases in the segmentation refinement process. Therefore, if the segmentation refinement fails, we can return to the existing coarse mask and consider other segmentation refinement methods. For this visualization, we used images publicly available on Wikimedia Commons for support-set. Query images are from PASCAL. We selected images where only one object was visible for each class. The support mask was obtained using the SAM segmentation method.
@@
How JFS works?: Most FSS networks [11, 15, 16, 20, 21, 22] rely on high quality support set information to find target objects in the query image (Fig. 1). Therefore, FSS performance deteriorates when the support mask information is inaccurate [20, 21]. If a refined mask accurately segments the target object, it provides more detailed texture information of the target to FSS model. Hence, we designed JFS. We found from our qualitative analysis (Fig. 2) that the JFS method was particularly effective in the following cases :
@@
(b) When the difference between the coarse and refined mask is clear, the presence of non-target objects in the refined mask confuses the FSS model, significantly dropping support mask prediction accuracy. Thus, it becomes evident whether the refined mask has failed or succeeded. Table 1 shows the JFS success rate rises with significant mIoU improvement, indicating JFS’s higher effectiveness when the refinement module makes greater improvement.
@@
Fig. 4: Failure cases. When the refined mask and coarse mask cover the same key features on the same target object, support mask prediction yields similar results. This is because FSS models are trained to find a target even with small key features from support-set.
5 Conclusion
We proposed JFS, a novel method for assessing the effectiveness of the segmentation refinement process. JFS can distinguish between successful and failed cases in SEPL segmentation refinement. The identification of successful refinements can be applied in various scenarios, leading to improved segmentation quality. Furthermore, JFS introduces a fresh perspective in the few-shot segmentation and segmentation refinement domains by repurposing FSS for a unique application. However, we recognize that JFS underperforms in certain instances, indicating the need for further development (Fig. 4). Future research could achieve better results by fine-tuning FSS models or adapting them specifically for the JFS framework.
6 Acknowledgment
This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NERSC DDR-ERCAP0033558.